OpenStudy (anonymous):

which gragh represents this equation 3x+7y=42

OpenStudy (cwrw238):

is there an illustration?

OpenStudy (anonymous):

yea here it is lol and thanks for trying to help

OpenStudy (anonymous):

If we find out where this equation, 3x+7y=42, crosses the y-axis and the x-axis, we can figure out which of the 4 graphs. Where does it cross the x-axis at? If we are crossing the x-axis on the line, then that means that y=0. 3x+7y=42 3x+7(0)=42 [plug in 0 for y] 3x+ 0 =42 [divide both sides by 3] x=14 the point (14,0) is on the line function of 3x+7y=42. Do any of the graphs have a line that goes through the point (14,0) ?